OpenStudy (jenniferjuice):

(x-3)(x-3) simplify.

OpenStudy (anonymous):

x * x + x * 3 -3 * x - 3 * 3 then solve that Does this make sense?

OpenStudy (jenniferjuice):

so... x^2 + 3x - 3x - 9 ?

OpenStudy (anonymous):

You can solve that even further. 3x - 3x. so x^2 - 9

HanAkoSolo (jamierox4ev3r):

once again you can use the FOIL method. do you remebber it?

OpenStudy (jenniferjuice):

..so i think i would combine the like terms

OpenStudy (anonymous):

(x-3)(x-3) x * x = x^2 x * -3 = -3x -3 * x = -3x -3 * -3 = +9 so: x^2 - 3x - 3x + 9 Then: x^2 -6x + 9 which is: (x -3)(x-3)

HanAkoSolo (jamierox4ev3r):

(x-3)(x-3) x^2-3x-3x+9

OpenStudy (jenniferjuice):

x^2 + 6x - 9....?

HanAkoSolo (jamierox4ev3r):

then your final answer is x^2-6x+9, which can be factored back into (x-3)(x-3), just like @some_someone just mapped out here

HanAkoSolo (jamierox4ev3r):

no the 9 is positive because two negatives equal a positive number

OpenStudy (anonymous):

a negative - a negative = a negative a negative * a negative = a positive

HanAkoSolo (jamierox4ev3r):

right! @some_someone a negative multiplied by a positive is negative

OpenStudy (anonymous):

lol yup Jamie, i was just pointing it out for Jennifer :)

OpenStudy (anonymous):

Oh, oops. I didn't see that other negative sign.

OpenStudy (jenniferjuice):

so i did it wrong...or?

HanAkoSolo (jamierox4ev3r):

x^2-6x+9 is the final answer? does that make sense @jenniferjuice ? if it doesn't what questions do you have about it

OpenStudy (anonymous):

Well Jenifer, you first said: x^2 + 6x - 9 But if was x^2 - 6x + 9

HanAkoSolo (jamierox4ev3r):

it was close but you have to be particular with your negatives

OpenStudy (jenniferjuice):

okay so.... x * x + x * 3 +3 * x - 3 * 3 then, x^2 + 3x + 3x - 9 then, x^2-6x+9

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x^2-9 = (x + 3)(x - 3)

OpenStudy (anonymous):

not quite @zzr0ck3r

OpenStudy (zzr0ck3r):

woops lol

OpenStudy (zzr0ck3r):

its been a long say I should take a break after that note:)